# Actor-Critic Model Predictive Control

### v5

@@ Abstract
An open research question in robotics is how to combine the benefits of model-free reinforcement learning (RL)—known for its strong task performance and flexibility in optimizing general reward formulations—with the robustness and online replanning capabilities of model predictive control (MPC). This paper provides an answer by introducing a new framework called Actor-Critic Model Predictive Control. The key idea is to embed a differentiable MPC within an actor-critic RL framework. The proposed approach leverages the short-term predictive optimization capabilities of MPC with the exploratory and end-to-end training properties of RL. The resulting policy effectively manages both short-term decisions through the MPC-based actor and long-term prediction via the critic network, unifying the benefits of both model-based control and end-to-end learning. We validate our method in both simulation and the real world with a quadcopter platform across various high-level tasks. We show that the proposed architecture can achieve real-time control performance, learn complex behaviors via trial and error, and retain the predictive properties of the MPC to better handle out of distribution behaviour.
@@ I Introduction
However, as tasks grow in complexity, model-based approaches necessitate substantial manual engineering, tailored to each specific task. This includes the careful crafting of the cost function and design of an appropriate planning strategy [10, 12]. Often, conservative assumptions about the task are made, leading to potentially sub-optimal task performance, for instance, in tasks where the dynamical system is taken to its limits [9, 10], or in applications that require discrete mode-switching [13]. Furthermore, the modular structure of model-based approaches may result in the progressive build-up of errors, accumulating in a cascading manner. This can compound inaccuracies, reinforce conservative estimations, and diminish the overall effectiveness of the system [14, 15, 16].
Figure 1: A block diagram of the approach. We combine the strength of actor-critic RL and the robustness of MPC by placing a differentiable MPC as the last layer of the actor policy. At deployment time the commands for the environment are drawn from solving an MPC, which leverages the dynamics of the system and finds the optimal solution given the current state.
Most recently, reinforcement-learning-based control has gained considerable traction, demonstrating exceptional performance in various domains, such as board games [17], video games [18], and drone racing both state-based [14, 19] and vision-based [20, 21]. Some of the most impressive achievements in robotics [14, 20, 22, 13] using reinforcement learning (RL) are even beyond the reach of existing model-based control methods. RL offers several advantages over MPC in robotics, notably in adaptability and flexibility [14]. RL can optimize policies directly from interactions with the environment, making it more flexible in defining the task goal. This flexibility and adaptability can lead to more scalable solutions, particularly in complex environments where MPC may struggle to provide optimal solutions [14].
@@
However, RL architectures are not without their own set of challenges [15, 23]. Training end-to-end without incorporating and leveraging prior knowledge, such as physics or dynamic models, results in need to learn everything from data. While the end-to-end paradigm is attractive, it demands a substantial amount of data and often lacks in terms of generalizability and robustness to out-of-distribution scenarios. This has resulted in hesitancy in applying end-to-end learned architectures to safety-critical applications and has fostered the development of approaches that advocate for the introduction of safety in learned pipelines [23, 24, 25].
@@
In this work, we introduce a new architecture called Actor-Critic Model Predictive Control to bridge the gap between Reinforcement Learning and Model Predictive Control. This architecture equips the agent with a differentiable MPC [26], located at the last layer of the actor network, as shown in Fig. 1, that provides the system with online replanning capabilities and allows the policy to predict and optimize the short-term consequences of its actions. Instead of relying in intermediate representations such as trajectories, we directly learn a map from observations to cost. Therefore, at deployment time, the control commands are drawn from solving an MPC, which leverages the system’s dynamics and finds the optimal solution given the current state. The differentiable MPC module, which incorporates a model of the system’s dynamics, provides the agent with prior knowledge even before any training data is received. The second component of our actor is the cost map, a deep neural network that encapsulates the dependencies between observations and the cost function of the MPC. In other words, while the MPC captures temporal variations inside its horizon, the neural cost module encodes the dependencies in relation to the observations. This architecture thereby incorporates two different time horizon scales: the MPC drives the short-term actions while the critic network manages the long-term ones. We demonstrate that our approach can tackle the agile flight problem with a highly non-linear quadrotor system, validated in both simulation and in real-world deployment.
@@ II Related Work
Several methods have been developed to learn cost functions or dynamic models for MPC [27, 28, 29, 30, 31, 32]. For example, in [29, 30], a policy search strategy is adopted that allows for learning the hyperparameters of a loss function for complex agile flight tasks. On the other hand, in [31, 32] they use Bayesian Optimization to tune the hyperparameters and dynamics of MPC controllers for different tasks such as car racing. However, these approaches use black-box optimization methods and do not exploit the gradient through the optimization problem, thus cannot leverage the full advantage of the prior knowledge embedded in the MPC.
Sampling-based MPC algorithms [33], are designed to handle intricate cost criteria and general nonlinear dynamics. This is accomplished by integrating neural networks for approximating system dynamics with the Model Predictive Path Integral (MPPI) control framework [33] for optimizing control in real-time. A vital aspect of sampling-based MPC is the generation of a large number of samples on-the-fly, often carried out in parallel using Graphics Processing Units (GPUs). Consequently, running sampling-based MPC on embedded systems can be both computationally demanding and memory intensive.
Alternatively, approaches leveraging differentiability through controllers have been on the rise. For example, for tuning linear controllers by getting the analytic gradients [34], for differentiating through an optimization problem for planning the trajectory for a legged robot [35], or for creating a differentiable prediction, planning and controller pipeline for autonomous vehicles [36]. On this same direction, MPC with differentiable optimization [26, 37, 38, 39] proposed to learn the cost or dynamics of a controller end-to-end. This approach is facilitated by analytically differentiating through the fixed point of a nonlinear iLQR solver [40]. Consequently, this method boasts substantial efficiency: it is less demanding in terms of computation and memory. However, all these approaches were only demonstrated in the context of imitation learning. While imitation learning is effective, its heavy reliance on expert demonstrations is a constraint. This dependence prevents exploration, potentially inhibiting its broader capabilities.
@@
We address this issue by leveraging the advantages of both differentiable MPC and model-free reinforcement learning. By equipping the actor with a differentiable MPC, our approach provides the agent with online replanning capabilities and with prior knowledge, which is a significant advantage over model-free RL, where the actor is a randomly initialized feedforward neural network. Unlike conventional MPC, our approach emphasizes robustness and adaptability, flexibly allowing for the optimization of intricate objectives through iterative exploration and refinement.
@@ III-A Preliminaries
Consider the discrete-time dynamic system with continuous state and input spaces, $\bm{x}_{k}\in\mathcal{X}$ and $\bm{u}_{k}\in\mathcal{U}$ respectively. Let us denote the time discretized evolution of the system $f:\mathcal{X}\times\mathcal{U}\mapsto\mathcal{X}$ such that $\bm{x}_{k+1}=f(\bm{x}_{k},\bm{u}_{k})$, where the sub-index $k$ is used to denote states and inputs at time $t_{k}$. The general Optimal Control Problem considers the task of finding a control policy $\pi(\bm{x})$, a map from the current state to the optimal input, $\pi:\mathcal{X}\mapsto\mathcal{U}$, such that the cost function $J:\mathcal{X}\mapsto\mathbb{R}^{+}$ is minimized:
$\pi(\bm{x})=$ $\operatorname*{argmin}_{u}$ $J(\bm{x})$
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bm{u}_{k}\in\mathcal{U}}$ (1)
where $k$ ranges from 00 to $N$ for $x_{k}$ and from 00 to $N-1$ for $\bm{u}_{k}$.
III-B General Quadratic MPC formulation
@@
Most MPC approaches need an explicit manual selection of a cost function that properly encodes the end task. For a standard tracking MPC this encoding is done through planning by finding a dynamically feasible reference trajectory that translates the task into suitable cost function coefficients for every time step. However, this approach presents two drawbacks: i) finding a dense, differentiable cost function can be difficult, and ii) even if this cost function is found, extra effort needs to be spent in fine tuning the parameters for real-world deployment. More generally, all receding horizon architectures such as MPC need to run in real-time when a deployment in the real world is desired. Because of this, the optimization problem is often approximated and converted from a non-linear optimization problem to a Quadratic Program (QP). Therefore, a more general cost function can be written as in Eq. (2).
$J_{Q}(\bm{x})=\sum_{k=0}^{N}\begin{bmatrix}x_{k}\\ u_{k}\end{bmatrix}^{T}Q_{k}\begin{bmatrix}x_{k}\\ u_{k}\end{bmatrix}+p_{k}\begin{bmatrix}x_{k}\\ u_{k}\end{bmatrix}$ (2)
In our paper, we propose to directly search for the matrix coefficients of Eq. (2). This way, by varying $Q_{k}$ and $p_{k}$, we are able to capture a larger family of problems, without suffering from the dependency on a feasible trajectory.
III-C Actor-Critic Reinforcement Learning
The Actor-Critic method is a widely-used approach in reinforcement learning (RL) that combines the advantages of both value-based and policy-based methods. It consists of two main components: the Actor and the Critic. The key idea is to simultaneously learn a state-value function $V_{\omega}(s)$ and learn a policy function $\pi_{\theta}$, where the value function (Critic) and policy (Actor) are parameterized by $\omega$ and $\theta$ separately. The policy is updated via the policy gradient theorem [41],
$\nabla_{\theta}J(\pi_{\theta})=\frac{1}{N}\sum_{i=1}^{N}\sum_{k=1}^{T}\nabla_{\theta}\log\pi_{\theta}(a_{k}^{i}|s_{k}^{i})A_{\omega}(s_{k}^{i},a_{k}^{i})$ (3)

### v6

@@ Abstract
An open research question in robotics is how to combine the benefits of model-free reinforcement learning (RL)—known for its strong task performance and flexibility in optimizing general reward formulations—with the robustness and online replanning capabilities of model predictive control (MPC). This paper provides an answer by introducing a new framework called Actor-Critic Model Predictive Control. The key idea is to embed a differentiable MPC within an actor-critic RL framework. This integration allows for short-term predictive optimization of control actions through MPC, while leveraging RL for end-to-end learning and exploration over longer horizons. Through various ablation studies, we expose the benefits of the proposed approach: it achieves better out-of-distribution behaviour, better robustness to changes in the dynamics and improved sample efficiency. Additionally, we conduct an empirical analysis that reveals a relationship between the critic’s learned value function and the cost function of the differentiable MPC, providing a deeper understanding of the interplay between the critic’s value and the MPC cost functions. Finally, we validate our method in the drone racing task in various tracks, in both simulation and the real world. Our method achieves the same superhuman performance as state-of-the-art model-free RL, showcasing speeds of up to 21 m/s. We show that the proposed architecture can achieve real-time control performance, learn complex behaviors via trial and error, and retain the predictive properties of the MPC to better handle out of distribution behavior.
@@ I Introduction
Figure 1: Top: A block diagram of an actor-critic reinforcement learning architecture with a Multilayer Percepton (MLP) Bottom: A block diagram of the proposed approach. We combine the strength of actor-critic RL and the robustness of MPC by placing a differentiable MPC as the last module of the actor policy. At deployment time the commands for the environment are drawn from solving an MPC, which leverages the dynamics of the system and finds the optimal solution given the current state. We show that the proposed approach achieves better out-of-distribution behaviour and better robustness to changes in the dynamics. We also show that the predictions of the differentiable MPC can be used to improve the learning of the value function and the sample efficiency.
@@
However, as tasks grow in complexity, model-based approaches necessitate substantial manual engineering, tailored to each specific task. This includes the careful crafting of the cost function, tuning of hyperparameters, and design of an appropriate planning strategy [10, 12]. Often, conservative assumptions about the task are made, leading to potentially sub-optimal task performance, for instance, in tasks where the dynamical system is taken to its limits [9, 10], or in applications that require discrete mode-switching [13]. Furthermore, the modular structure of model-based approaches may result in the progressive build-up of errors, accumulating in a cascading manner. This can compound inaccuracies, reinforce conservative estimations, and diminish the overall effectiveness of the system [14, 15, 16].
Most recently, reinforcement-learning-based control has gained considerable traction, demonstrating exceptional performance in various domains, such as board games [17], atari games [18], complex video games [19, 20], and drone racing, both state-based [14, 21] and vision-based [22, 23]. In particular, the results in [14] demonstrate that model-free RL can achieve superior performance compared to optimal control techniques in drone racing. This success is largely due to RL’s ability to directly optimize a non-differentiable, nonlinear, sparse objective, eliminating the need for proxy objectives in the form of a predefined reference time trajectory or continuous 3D path. Additionally, RL learns a control policy through offline optimization, enabling efficient real-time control during deployment. Unlike trajectory optimization, it learns a feedback controller for real-time adaptability and directly optimizes task objectives, bypassing intermediate representations. Finally, RL can leverage domain randomization to cope with model uncertainty, allowing the discovery of more robust control responses. Unlike standard optimal control methods, RL optimizes policies directly through interactions with the environment. This allows RL to handle sparse reward functions (e.g., minimize number of crashes, minimize total energy consumption, minimize lap times, etc.). Additionally, RL can enhance robustness and prevent overfitting through domain randomization, which involves training the agent across a range of simulated environments with diverse settings. This flexibility and adaptability can lead to more scalable solutions, particularly in complex environments where MPC may struggle to provide optimal solutions [14].
@@
However, RL architectures are not without their own set of challenges [15, 24]. Learning a model-free, end-to-end policy without explicitly leveraging prior knowledge in the training process, such as physics or dynamic models, results in the need to learn everything from data – often resulting in millions of interactions within a simulator. While the end-to-end paradigm is attractive, it often lacks in terms of generalizability and robustness to out-of-distribution scenarios. This has resulted in hesitancy in applying end-to-end learned architectures to safety-critical applications and has fostered the development of approaches that advocate for the introduction of safety in learned pipelines [24, 25, 26].
This article is an extension of a previous work [27]. In [27], we introduced a new architecture called Actor-Critic Model Predictive Control to bridge the gap between Reinforcement Learning and Model Predictive Control. This architecture equips the agent with a differentiable MPC [28], placed as the last module of the actor network, as shown in Fig. 1, that provides the system with online replanning capabilities and allows the policy to predict and optimize the short-term consequences of its actions.
@@
Instead of relying on intermediate representations such as trajectories, we directly learn a map from observations to the cost function. Therefore, at deployment time, the control commands are drawn from solving an MPC, which leverages the system’s dynamics and finds the optimal solution given the current state. The differentiable MPC module, which incorporates a model of the system’s dynamics, provides the agent with prior knowledge even before any training data is received. The second component of our actor is the cost map, a deep neural network that encapsulates the dependencies between observations and the cost function of the MPC. In other words, while the differentiable MPC captures temporal variations inside its horizon, the neural cost module encodes the dependencies in relation to the observations. This architecture thereby incorporates two different time horizon scales: the MPC drives the short-term actions while the critic network manages the long-term ones. This represents a significant advantage over vanilla actor-critic RL, where the actor is typically a randomly initialized feedforward neural network with no domain-specific structure nor priors.
In [27], we demonstrated that the AC-MPC architecture exhibits more stable behavior in out-of-distribution scenarios, highlighting its robustness and generalizability compared to its neural-network-only counterpart. Lastly, in [27], we showcased the applicability of the proposed approach in real-world scenarios using an agile quadrotor platform, demonstrating the feasibility and effectiveness of the AC-MPC architecture in dynamic and challenging environments.
This paper extends [27] as follows:
  • We propose an enhancement to the AC-MPC algorithm by incorporating predictions into the critic’s training process. This improvement leverages the prior knowledge of the system through the MPC predictions and incorporates them in the learning process the value function, resulting in improved sample efficiency during training.
  • We conduct an empirical analysis that reveals a relationship between the Hessian of the learned value function and the learned cost functions of the differentiable MPC. This finding provides a deeper understanding of the interplay between the RL value and the MPC cost functions within the AC-MPC framework, offering valuable insights for future research.
  • We show that after optimizing hyperparameters, specifically exploration, AC-MPC is able to leverage the prior knowledge included in the dynamics of the system and achieve better training performance than AC-MLP (conventional PPO).
  • We show that our approach is more robust to dynamic parameter changes in comparison with a baseline, keeping high success rates for variations in mass, in XY inertia, and body rate limits. This improved robustness is attributed to the differentiable MPC’s access to the dynamics model.
  • We demonstrate through both simulation and real-world experiments that the proposed AC-MPC framework achieves performance on par with state-of-the-art model-free reinforcement learning approaches in the drone racing task, attaining speeds of up to 21 m/s on a quadrotor platform.
Overall, the AC-MPC architecture not only advances the integration of RL and MPC but also offers practical improvements in out of distribution behaviour, training efficiency, and real-world applicability, contributing to the development of more robust and efficient control systems.
@@ II Related Work
Model Predictive Control (MPC) in robotics. MPC has had an enormous success in robotic applications over the last decades [29, 30, 31, 32, 9, 10, 11, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46]. However, often the cost function needs to be crafted by hand by an expert, and the hyperparameters need to be tuned at deployment time [47, 48]. Additionally, in most MPC setups [10, 49, 50, 51, 52, 7, 53, 54, 9], the high-level task is first converted into a reference trajectory (planning) and then tracked by a controller (control). This decomposition of the problem into these distinct layers is greatly favored by the MPC methodology, largely due to the differentiability and continuity requirement of optimal control’s cost functions. The cost function is shaped to achieve accurate trajectory tracking and is decoupled from (and usually unrelated to) the high-level task objective. As a result, the hierarchical separation of the information between two components leads to systems that can become erratic in the presence of unmodeled dynamics. In practice, a series of conservative assumptions or approximations are required to counteract model mismatches and maintain controllability, resulting in systems that are no longer optimal [14, 15, 16].
Another promising model-based direction for addressing complex control challenges is sampling-based MPC algorithms [55, 56, 57, 58], which are designed to handle intricate, non-differentiable cost functions and general nonlinear dynamics. These algorithms integrate system dynamics – either known or learned – into the Model Predictive Path Integral (MPPI) control framework, optimizing control in real-time. A key feature of sampling-based MPC is the ability to generate a large number of control samples on-the-fly, often leveraging parallel computation via Graphics Processing Units (GPUs). However, implementing sampling-based MPC on embedded systems poses significant challenges, as it tends to be both computationally demanding and memory-intensive. Reinforcement Learning (RL) in robotics. RL has risen as an attractive alternative to conventional controller design, achieving impressive performance that goes beyond model-based control in a variety of domains [17, 59, 19, 60, 61, 62]. RL optimizes a controller using sampled data and can handle nonlinear, nonconvex, non-differentiable and even sparse objectives, manifesting great flexibility in controller design. Compared to model-based optimal control, RL has a number of key advantages: most importantly, RL can optimize the performance objective of the task directly, removing the need for explicit intermediate representations, such as trajectories. Moreover, RL can solve complex tasks from raw, high-dimensional sensory input, without the need of a specific metric state [14]. However, RL is still far from reaching the level of robustness and generalisability of model-based control approaches when deployed in the real world, due to its brittleness when deployed in situations outside of the training distribution, and its lack of guarantees [24]. Model-free methods generally use black-box optimization methods and do not exploit the first order gradient through the dynamics, thus cannot leverage the full advantage of the prior knowledge.
@@
Combinations of MPC and learning. Several methods have been developed to learn cost functions for MPC [63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77], dynamics models for MPC [40, 78, 79, 80, 81, 82], or both simultaneously [83, 42, 28, 84, 85, 86, 87, 88, 89]. For example, in [64, 65], a policy search strategy is adopted that allows for learning the hyperparameters of a loss function for complex agile flight tasks. On the other hand, the works in [40, 63] use Bayesian Optimization to tune the hyperparameters and dynamics of MPC controllers for different tasks such as car racing. Similarly, in [72] the authors propose a new method called Deep Model Predictive Optimization (DMPO) that learns the update rule of an MPPI controller using RL. They evaluate their algorihtm on a real quadrotor platform and outperform state-of-the-art MPPI algorithms.
Recent model-based RL approaches integrate MPC-like components either during policy learning [90, 91] or by using explicit MPC-based policies [92, 93, 83, 94]. Many of these methods learn a dynamics model and use it to perform MPC at inference time. The latter step is either done via random shooting [92], the cross-entropy method [93, 95], or a combination of these methods with a learned critic [96, 83].
Alternatively, approaches leveraging differentiability through optimizers have been on the rise. For example, for tuning linear controllers by getting the analytic gradients [97], for differentiating through an optimization problem for planning the trajectory for a legged robot [98], or for creating a differentiable prediction, planning and controller pipeline for autonomous vehicles [99], or for moving horizon estimation [100].
On this same direction, MPC with differentiable optimization [28, 101, 102, 103] proposed to learn the cost or dynamics of a controller end-to-end. In particular, the authors in [28] were able to recover the tuning hyperparameters of an MPC via imitation learning for non-linear, low-dimensional dynamics – such as cartpole and inverted pendulum – by backpropagating through the MPC itself. Later, in [104], the authors augment the cost function of a nominal MPC with a learned cost that uses the gradient through the optimizer for the task of navigating around humans. To train this learned cost, they also learn from demonstrations. Until [27], all these approaches were only demonstrated in the context of imitation learning. While imitation learning is effective, its heavy reliance on expert demonstrations is a burden. This dependence prevents exploration, potentially inhibiting its broader capabilities.
@@
In [27], we address this issue by leveraging the advantages of both differentiable MPC and model-free reinforcement learning. By equipping the actor with a differentiable MPC, our approach provides the agent with online replanning capabilities and with prior knowledge, which is a significant advantage over model-free RL, where the actor is a randomly initialized feedforward neural network. Unlike conventional MPC, our approach emphasizes robustness and adaptability, flexibly allowing for the optimization of intricate objectives through iterative exploration and refinement.
@@ III Methodology
In this section we first present preliminaries about both Optimal Control and Reinforcement Learning, and then introduce our method.
@@ III-A Preliminaries
Consider the discrete-time dynamic system with continuous state and input spaces, $\bm{x}_{k}\in\mathcal{X}$ and $\bm{u}_{k}\in\mathcal{U}$ respectively. Let us denote the time discretized evolution of the system $f:\mathcal{X}\times\mathcal{U}\mapsto\mathcal{X}$ such that $\bm{x}_{k+1}=f(\bm{x}_{k},\bm{u}_{k})$, where the sub-index $k$ is used to denote states and inputs at time $t_{k}$. The general Optimal Control Problem considers the task of finding a control policy $\pi(\bm{x})$, a map from the current state to the optimal input, $\pi:\mathcal{X}\mapsto\mathcal{U}$, such that the cost function $J:\mathcal{X}\mapsto\mathbb{R}^{+}$ is minimized,
$\pi(\bm{x})=$ $\operatorname*{argmin}_{\bm{u}}$ $J(\bm{x})$
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bm{u}_{k}\in\mathcal{U}},$ (1)
where $k$ ranges from 00 to $N$ for $x_{k}$ and from 00 to $N-1$ for $\bm{u}_{k}$. In the following sections, we will go through different ways in which $J(\bm{x})$ can be chosen.
III-B Tracking Model Predictive Control
For tracking Model Predictive Control approaches [105, 106, 107], the objective $J(\bm{x})$ is to minimize a quadratic penalty on the error between the predicted states and inputs, and a given dynamically feasible reference $\bm{x}_{k,ref}$ and $\bm{u}_{k,ref}$. Consequently, the cost function $J(\bm{x})$ in problem (1) is substituted by:
$J_{MPC}(\bm{x})=\sum_{k=0}^{N-1}\|\Delta\bm{x}_{k}\|_{\bm{Q}}^{2}+\|\Delta\bm{u}_{k}\|_{\bm{R}}^{2}+\|\Delta\bm{x}_{N}\|_{\bm{P}}^{2},$ (2)
where $\Delta\bm{x}_{k}=\bm{x}_{k}-\bm{x}_{k,ref}$, $\Delta\bm{u}_{k}=\bm{u}_{k}-\bm{u}_{k,ref}$, and where $\bm{Q}\succeq 0$, $\bm{R}\succ 0$ and $\bm{P}\succeq 0$ are the state, input and final state weighting matrices. The norms of the form $\|\cdot\|_{A}^{2}$ represent the weighted Euclidean inner product $\|\bm{v}\|^{2}_{A}=\bm{v}^{T}A\bm{v}$.
This formulation relies on the fact that a feasible reference $\bm{x}_{k,ref}$, $\bm{u}_{k,ref}$ is accessible for the future time horizon $N$. Searching for these is often referred to as planning, which, depending on the application, can be both computationally intensive and complex, particularly in scenarios involving cluttered or dynamic environments [9, 108, 109]. In many cases, solving the planning problem becomes the bottleneck in real-time applications due to high computational requirements and the need for precise system models.
However, the tracking MPC formulation also relies on the assumption that a trajectory represents a well-posed, quadratic proxy objective for solving the end task. While this assumption holds for many practical use cases, it becomes limiting in situations where the system operates near its performance limits, such as aggressive maneuvers or high-speed tasks, where the trajectory may no longer serve as an effective proxy. In such cases, the tracking objective can result in suboptimal performance or even task failure [14].
III-C Economic Model Predictive Control
On the other hand, economic MPC offers an alternative by directly optimizing the task objective. Unlike traditional tracking MPC, economic MPC allows for a more flexible design of the cost function, which can vary depending on the specific task [110, 111].
$J_{EMPC}(\bm{x})=\sum_{k=0}^{N-1}l(x_{k},u_{k}),$ (3)
where $l(x_{k},u_{k})$ is directly the stage cost. This stage cost is generally designed to optimize sparse metrics, such as energy consumption, efficiency, or success rate.
While economic MPC offers advantages in terms of flexibility with respect to task design, it also presents several challenges that must be addressed for successful implementation. One of the primary caveats is that the stage cost $l(x_{k},u_{k})$ in most tasks tends to be sparse, non-differentiable and non-smooth, which complicates the formulation of Eq. (3). For example, in tasks with discontinuous behaviour – e.g., switching between different modes, dealing with contact forces, or dealing with agent crashes – it is generally difficult to find a cost function that is differentiable.
III-D General Quadratic MPC formulation
In most MPC approaches there is a need of i) an explicit manual selection of a differentiable cost function that properly encodes the end task, and ii) hyperparameter tuning. As mentioned in section III-B, in the case of a standard tracking MPC this encoding is done through planning, by finding a dynamically feasible reference trajectory that translates the task into suitable cost function coefficients for every time step. For economic MPC (section III-C), recent research has been going in the direction of using data to approximate the cost function [112], or making it computationally tractable by transforming it to a quadratic program using the Hessian [113, 114].
@@ III-D General Quadratic MPC formulation
However, these approaches present two main drawbacks: i) hand-crafting a dense, differentiable cost function can be difficult for a general task, and ii) even if this cost function is found, extra effort needs to be spent in fine tuning the hyperparameters for real-world deployment. More generally, optimization-based architectures such as MPC need to run in real-time when a deployment in the real world is desired. As a result, the underlying optimization problem is often simplified by approximating the original nonlinear formulation and converting it into a Quadratic Program (QP) suitable for the real-time iteration (RTI) [115, 116, 117] scheme to ensure computational tractability. A general quadratic cost function can be written as
$J_{QP}(\bm{x})=\sum_{k=0}^{N}\begin{bmatrix}x_{k}\\ u_{k}\end{bmatrix}^{T}Q_{k}\begin{bmatrix}x_{k}\\ u_{k}\end{bmatrix}+p_{k}\begin{bmatrix}x_{k}\\ u_{k}\end{bmatrix}.$ (4)
In this paper, we propose to directly search for the matrix coefficients of Eq. (4) – $Q_{k}$ and $p_{k}$ – using reinforcement learning and differentiable MPC [28]. This way we are able to encode the task as a general, non-differentiable reward function while at the same time maintaining computational tractability.
III-E Actor-Critic Reinforcement Learning
Reinforcement learning problems are often framed within the Markov Decision Process (MDP) formalism, which provides a structured approach for modeling decision-making problems. An MDP is defined by the tuple $(\mathcal{S},\mathcal{A},P,R,\gamma)$, where $\mathcal{S}$ is the set of possible states, $\mathcal{A}$ is the set of possible actions, $P(s_{k+1}|s_{k},a_{k})$ is the state transition probability, $R(s_{k},a_{k})$ is the reward function, and $\gamma\in[0,1)$ is the discount factor. The objective in an MDP is to find a policy $\pi$ that maximizes the expected cumulative reward over time. Within the context of Problem (1), and when the state transition function is deterministic, it is equivalent to the system dynamics $f(x_{k},u_{k})$, where in this case the observation $s_{k}$ is directly the state $x_{k}$ and the action $a_{k}$ is directly the control input $u_{k}$.
To solve such problems, a common approach in Reinforcement Learning is to optimize directly the policy $\pi_{\theta}$ by using the policy gradient. The policy $\pi_{\theta}$ is parameterized by $\theta$, typically weights of a neural network. The goal is to adjust these parameters to maximize the expected return, defined in the infinite-horizon case as:
$R(\tau)=\sum_{k=0}^{\infty}\gamma^{k}r(s_{k},a_{k}),$ (5)
where $\tau$ denotes a trajectory, $r(s_{k},a_{k})$ the reward function, and $\gamma\in[0,1)$ is a discount factor that discounts future rewards. Unlike real-time optimization in methods like MPC, RL can directly optimize a non-differentiable, sparse reward. Additionally, policy gradient optimization is typically performed offline through interactions with simulation. Once trained, the policy $\pi_{\theta}$ enables the computation of control signals by simply evaluating the function $a^{\ast}=\pi_{\theta}(s_{k})$ at each time step, reducing the computational complexity during deployment. For Actor-Critic reinforcement learning, the key idea is to simultaneously learn a state-value function $V_{\omega}(s)$ and learn a policy function $\pi_{\theta}$, where the value function (Critic) and policy (Actor) are parameterized by $\omega$ and $\theta$ separately. The policy is updated via the policy gradient theorem [118],
$\nabla_{\theta}J(\pi_{\theta})=\frac{1}{N}\sum_{i=1}^{N}\sum_{k=1}^{T}\nabla_{\theta}\log\pi_{\theta}(a_{k}^{i}|s_{k}^{i})A_{\omega}(s_{k}^{i},a_{k}^{i}),$ (6)